(1) If inflation in Sweden is projected at 5% annually for the next 5 years and 12% in Turkey. If Lira/Krona is 205.56, calculate the exact relative PPP value of the spot rate 5 years from now?

(2) Assume that the spot exchange rate (SF/$) is currently 1.8960. U.S treasury securities with 3 years to maturity currently yield 6.68% per year, while the interest rate on comparable Swiss securities is 7.44%. Assuming annual compounding, what is the expected spot exchange rate SF/$ in 3 years if IFE holds exactly?
Answer #1
1] Relative PPP value = 205.56*(1.12/1.05)^5 = 283.8456
2] Expected spot rate = 1.8960*(1.0744/1.0668)^3 = 1.9368